The size P of a certain insect population at time t​ (in days) obeys the function P(t)=900e^0.03t
​(a) Determine the number of insects at t=0 days.
​(b) What is the growth rate of the insect​ population?
​(c) What is the population after 10​ days?
​(d) When will the insect population reach 1440?
​(e) When will the insect population​ double?

Found 2 solutions by Boreal, Theo:
Answer by Boreal(15235)

growth rate is 3% the 0.03 in the e^rt, where r is the rate of growth.
in 10 days the population is 900*e^0.3=1214.87 or 1215
reaches 1440 when 1440=900*e^0.03t
1.6=e^0.03 t
ln both sides
0.47=0.03t
t=15 2/3 days
doubles in 24 days using the rule of 72, but that is empirical
2=e^0.03t
ln2=0.03t
0.693=0.03t
t=23.1 days
Some use the rule of 70, which would give 23.3 days and is a better estimate.

Answer by Theo(13342)

p(t) = 900 * e^(.03*t)
p(t) = number of insects at t days.
t = number of days.

at t = 0, the population is equal to 900 * e^(.03*0) = 900 * e^0 = 900 * 1 = 900.

at t = 10, the population is equal to 900 * e^(.03*10) = 900 * e^(.3) = 1214.872927.
since you can't have part of an ant, round the answer to the nearest integer = 1215.

the formula for the number of days when the population will be equal to 1414 is 1440 = 900 * e^(.03*t)
divide both sides of this equation by 900 to get:
1440/900 = e^(.03*t)
take the natural log of both sides of the equation to get ln(1440/900) = ln(e^(.03*t))
since ln(e^x) = x*ln(e), this becomes ln(1440/900) = .03*t * ln(e)
since ln(e) = 1, this becomes ln(1440/900) = .03*t
solve for t to get t = ln(1440/900) / .03 = 15.66678764 days.
to confirm this is correct, replace t with 15.0594361 in the original equation to get p(t) = 900 * e^(.03 * 15.66678764) = 1440.

the insect population will double when it becomes 1800.

the formula for the number of days when the population will be equal to 1800 is 1800 = 900 * e^(.03*t)
divide both sides of this equation by 900 to get:
1800/900 = e^(.03*t)
take the natural log of both sides of the equation to get ln(1800/900) = ln(e^(.03*t))
since ln(e^x) = x*ln(e), this becomes ln(1800/900) = .03*t * ln(e)
since ln(e) = 1, this becomes ln(1800/900) = .03*t
solve for t to get t = ln(1800/900) / .03 = 23.10490602 days.
to confirm this is correct, replace t with 23.10490602 in the original equation to get p(t) = 900 * e^(.03 * 23.10490602) = 1800.

to find the daily growth rate, do the following.
start with 900.
multiply it by e^(.03 * 1 to get 927.4090806.
divide that by 900 to get 1.030454534.
subtract 1 from that to get .030454534 = 3.0454534%.
that's the daily growth rate.

to confirm, use that in the discrete growth rate formula of f = p * ( 1 + r) ^ n
f is the future value
p is the present value
r is the growth rate per time period.
n is the number of time periods.

for example:
the population in 10 days will be equal to 900 * (1 + .030454534) ^ 10 = 1214.872927.
this is the same result we got before with p(t) = 900 * e^(.03 * 10)
p(t) means the same as f.
they both represent the future value of the insect population.
p(t) and f tell you what the future value is.
p tells you what the present value is.

the discrete growth rate formula of f = p * (1 + .030454534) ^ n will give you the same answer as the continuous growth rate formula of p(t) = p * e^(.03 * t)